OpenStudy (anonymous):

1. Solve the expression below. (7 × 8) - 2 = ???

OpenStudy (anonymous):

Whats 7x 8?

OpenStudy (anonymous):

56- 2 = 54

OpenStudy (anonymous):

Multiply 7 times 8 first. (7 × 8) - 2 = 56 - 2 Then do the subtraction: 56 - 2 = 54

OpenStudy (anonymous):

I have no clue what melissaaaa did, but what ever it was, she's wrong.

OpenStudy (anonymous):

Yes @tim.gibbs

OpenStudy (anonymous):

yeah i just looked at it wrong , i corrected myself as well.

OpenStudy (anonymous):

You can use BEDMAS Brackets/Parentheses Exponents Division Multiplication Addition Subtraction In that order.

OpenStudy (anonymous):

2. Solve the expression below. 30 - (3 × 9) = ???

OpenStudy (anonymous):

Brackets first

OpenStudy (anonymous):

27

OpenStudy (anonymous):

Yes now subtract

OpenStudy (anonymous):

33

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

. Solve the expression below. (6 × 6) - 4 = ???

OpenStudy (anonymous):

Brackets :P

OpenStudy (zzr0ck3r):

@melissaaaa maybe you should learn from this situation on how to treat others, i.e. not laugh at them when you think they are wrong, not only is it rude, but sometimes you might be the wrong person.

OpenStudy (anonymous):

36

OpenStudy (anonymous):

32

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

Solve the expression below. 4 + (16 ÷ 2) = ???

OpenStudy (anonymous):

Brackets :)

OpenStudy (anonymous):

Remember what I showed you earlier? BEDMAS, follow that top to bottom.

OpenStudy (anonymous):

Or left to right lol

OpenStudy (anonymous):

8. Solve the expression below. 63 - (8 × 7) = ???

OpenStudy (anonymous):

One at a time, lets do 4+(16/2) first

OpenStudy (anonymous):

What's 16/ 2?

OpenStudy (anonymous):

63-56

OpenStudy (anonymous):

Yes, for the second one, but lets do 4+(16/2)

OpenStudy (anonymous):

What's 16 divided by 2?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Good job!